A Boeing 747 "Jumbo Jet" has a length of 65.0 m. The runway on which the plane lands intersects another runway. The width of the intersection is 24.2 m. The plane decelerates through the intersection at a rate of 5.55 m/s2 and clears it with a final speed of 44.4 m/s. How much time is needed for the plane to clear the intersection?
Answer #1

let

d = 24.2 m

a = -5.55 m/s^2

v2 = 44.4 m/s

let v1 is the speed of plane when starts crossing the intersection.


use,

v2^2 - v1^2 = 2*a*d

v1^2 = v2^2 - 2*a*d

v1 = sqrt(44.4^2 - 2*(-5.55)*24.2)

= 47.3 m/s

time taken, t = (v2 - v1)/a

= (44.4 - 47.3)/(-5.55)

= 0.522 s <<<<<<<<---------------Answer